# Coexistence of Anderson Localization and Quantum Scarring in Two Dimensions

### v1

@@ Abstract
We study finite two-dimensional disordered systems with periodic confinement. At low energies, eigenstates exhibit strong Anderson localization, while at higher energies a subset of states forms variational scars with anisotropic intensity patterns that violate random wave expectations. Scaling theory predicts that all states localize in two dimensions, yet energy-dependent localization lengths and finite system size allow these regimes to coexist. We demonstrate that this coexistence produces distinct, robust signatures in both spatial intensity patterns and spectral statistics that are directly observable in mesoscopic electronic, photonic, and cold atom systems.
Introduction– In disordered quantum systems, the breakdown of ergodicity can manifest in several distinct forms, from Anderson localization (AL) in non-interacting settings to weakly non-ergodic structures that defy random‑wave expectations [DeTomasi2020Multifractality, Altshuler2016Nonergodic, Das2025Emergent, Xu2023NEE]. Anderson’s original insight is that quenched disorder can localize single-particle eigenstates and suppress transport. This principle establish an essential tenet of mesoscopic physics [Anderson1958, Abrahams1979, Billy2008, Roati2008]. In parallel, deviant wavefunctions that concentrate along classical periodic orbits (POs), known as quantum scars [Heller_book], offer a different route to non-ergodic behavior within single-particle systems. Such scarred states are reported in billiard-like platforms [Heller_phys.rev.lett_53_1515_1984, Noeckel1997, Bogomolny2012] as well as in quantum wells [Luukko_sci.rep_6_37656_2016, keski-rahkonen_phys.rev.lett_123_214101_2019, Selinummi2024, Chalangari2025, keskirahkonen_phys.rev.e_112_L012201_2025], and recently, scarred eigenstates were directly visualized in a graphene-based quantum well [Ge_nature_635_841_2024]. Many-body analogs of these scars have also been uncovered in constrained lattice models [Turner2018, Serbyn2021]. These phenomena are nowadays routinely probed across platforms with high control, such as two‑dimensional (2D) electron gases, disordered photonic media, waveguide arrays, and ultracold atoms  [Schwartz2007, Lahini2008, Kondov2011, directVis_Ge2024, PhysRevB103104201, White2020].
@@
Despite the progress on these frontiers, the manner in which finite-size effects and continuum disorder conspire to yield coexisting non-ergodic regimes across the spectrum in realistic systems remains elusive. To address this issue, we here investigate a finite 2D continuum Hamiltonian with prototypical disorder of Gaussian potential bumps serving as a proxy for local impurities. Our analysis reveals regimes that separate across the energy spectrum: (i) conventional AL eigenstates dominating the low-energy sector, (ii) genuinely ergodic delocalized states following random-matrix-theory statistics in the sense of the Bohigas-Giannoni-Schmit conjecture [Bohigas_phys.rev.lett_52_1_1984], while (iii) a set of high-energy scarred states exhibiting intensity patterns that deviate from random-wave predictions of the Berry conjecture [Berry_j.phys.a_10_2083_1977]. Representative eigenstates of these classes are shown in Fig. 1 (upper panel and in Supplemental Material (SM) S1), illustrating a spectrum of behaviors ranging from AL [Fig. 1(a)], to ergodic delocalization [Fig. 1(b)] and varying degrees of scarring [Figs. 1(c & d)]. These regimes are distinguished through an analysis based on complementary observables, including spatial profiles, inverse participation ratios (IPR), multifractal spectra, and deviations from universal level statistics.
Model Setting– We consider a 2D potential profile $V_{\mathrm{ext}}(\mathbf{r})$, confined by Dirichlet hard-wall boundaries to a region of size $L\times L$, where $L$ denotes the number of quantum-wells along each direction, together with Gaussian impurities $V_{\text{imp}}(\mathbf{r})$. By utilizing the imaginary-time propagation method [LuukkoP.J.J.2013Itpc], we obtain the lowest $\sim 4\times 10^{3}-10^{4}$ eigenstates of the time-independent Schrödinger equation $H\psi_{n}(\mathbf{r})=E_{n}\psi_{n}(\mathbf{r})$ with
@@
with amplitude $A_{i}$, width $\sigma_{i}$, and position $\mathbf{r}_{i}$, which is a standard approach for modeling imperfections or impurities in quantum wells [hirose_phys.Rev.B_63_075301_2001, hirose_phys.rev.B_65_193305_2002]. The bump positions $\mathbf{r}_{i}$ are sampled independently from a uniform distribution, yielding an average density $\rho$ per unit area. The average impurity width is fixed in all simulations at $\langle\sigma_{i}\rangle=0.2$. The dimensionless ratio $\langle A\rangle/V_{0}$, where $\langle A\rangle$ is the average bump amplitude, characterizes the overall disorder strength.
@@
Figure 1 (lower panel) illustrates a typical localization–delocalization crossover in the disorder–energy plane for a system with fixed size $L$ and impurity density $\rho$. The colormap shows $\log_{10}(\mathrm{IPR}_{2})$, and the gray vertical dashed-line marks the scaled well-depth energy $V_{0}$. At energies below and near the well-depth, $E_{n}\lesssim V_{0}$, eigenstates are dominated by potential confinement (see a representative example in Fig.1(a)), and increasing disorder strength $\langle A\rangle/V_{0}$ disrupts inter-well resonances. This results in a broadening of the energy window populated by strongly AL states. At energies $E_{n}\gtrsim V_{0}$, the IPR landscape exhibits two quantitatively distinct regimes. One consists of very low IPR eigenstates in which scarring is essentially absent [see Fig. 1(b)]. For these states the localization length exceeds the system size, and de Broglie wavelength $\lambda(E)\sim 2\pi/\sqrt{2(E-\langle V\rangle)}$ $\langle V\rangle$ denotes the expectation value of total potential energy] is still not small enough to resolve the underlying periodic structure.
Upon further increasing the energy, de Broglie wavelength becomes sufficiently short to resolve the periodic structure of the confinement. In this regime, scarred states emerge as anisotropic outliers embedded within the delocalized background, corresponding to variational scarring. As the disorder strength increases, both the low- and intermediate- energy windows [i.e., Fig. 1(a) and (b)] broaden, pushing the scar-bearing region to progressively higher energies. In strongly disordered cases, the remaining scars are weakened and occupy a diminishing fraction of the accessible spectrum, consistent with previous observations [KeskiRahkonen2019b, simo].
@@
As discussed in Fig. 1, at energies $E_{n}\gtrsim V_{0}$, the spectrum contains both delocalized and scarred eigenstates. In 2D, the mean spacing between adjacent energy levels scales as $\sim L^{-2}$ [Girvin_Yang_2019]. Consequently, a fixed set of $10^{4}$ computed eigenstates spans a progressively narrower absolute energy window as the system size increases (see Fig. S3 in SM). Therefore, larger systems require increasingly greater number of eigenstates to reach energies at which the de Broglie wavelength becomes short enough to resolve the underlying geometry. The fixed sampled spectrum is thus dominated by delocalized states, while the scar-bearing window are pushed to higher relative energies.
The coexistence of AL, delocalized, and scarred states within the same sample can thus be explained by the interplay between the energy-dependent $\xi(E)$ and the finite system size. We find that scarred states emerge when (i) $\xi(E)$ becomes comparable to or exceeds $L$, and (ii) de Broglie wavelength becomes sufficiently small to allow eigenfunctions to resolve the underlying PO structure imposed by the bounded geometry. In this regime, the system admits families of unstable and marginally stable POs embedded within the chaotic sea, and interference along these trajectories imprints the characteristic directional, scarred patterns onto the corresponding eigenstates [kaplan_ann.phys_264_171_1998, Heller_book].
@@
Figure 3: Level-spacing statistics. (a) Distribution of consecutive level-spacing ratios for disordered systems of sizes $L=4$ and $L=13$. The corresponding symmetrized gap-ratio averages are $\langle\tilde{s}\rangle\approx 0.389$ and $0.524$, respectively. (b) Comparison of weak and strong disorder strengths, $\langle A\rangle/V_{0}=0.1$ and $2$, for $L=5$, yielding $\langle\tilde{s}\rangle\approx 0.388$ and $0.525$, respectively.
@@
Figure S3: (Left) $\text{IPR}_{2}$ versus energy comparing clean and disordered systems, and (right) $\text{IPR}_{2}$ versus energy for different system sizes, showing progressively narrower absolute energy interval as the system size increases .

### v2

@@ Abstract
We study finite two dimensional disordered systems with periodic confinement. At low energies, eigenstates exhibit strong Anderson localization, while at higher energies a subset of states forms variational scars with anisotropic intensity patterns that violate random wave expectations. Scaling theory predicts that all states localize for two dimensions in the thermodynamic limit, yet energy dependent localization length and finite system size allow these regimes to coexist. We demonstrate that this coexistence produces distinct, robust signatures in both spatial intensity patterns and spectral statistics that are directly observable in mesoscopic electronic, photonic, and cold atom systems.
Introduction– In disordered quantum systems, the breakdown of ergodicity can manifest in several distinct forms, from Anderson localization (AL) in non-interacting settings to weakly non-ergodic structures that defy random‑wave expectations [DeTomasi2020Multifractality, Altshuler2016Nonergodic, Das2025Emergent, Xu2023NEE]. Anderson’s original insight was that quenched disorder localizes single-particle eigenstates in one and two dimensions, and suppresses transport. This principle establish an essential tenet of mesoscopic physics [Anderson1958, Abrahams1979, Billy2008, Roati2008]. In parallel, deviant wavefunctions that concentrate along classical periodic orbits (POs), known as quantum scars [Heller_book], offer a different route to non-ergodic behavior within single-particle systems. Such scarred states are reported in billiard-like platforms [Heller_phys.rev.lett_53_1515_1984, Noeckel1997, Bogomolny2012] as well as in quantum wells [Luukko_sci.rep_6_37656_2016, keski-rahkonen_phys.rev.lett_123_214101_2019, Selinummi2024, Chalangari2025, keskirahkonen_phys.rev.e_112_L012201_2025], with recent direct visualization in a graphene-based quantum well [Ge_nature_635_841_2024]. Many-body analogs of these scars have also been uncovered in constrained lattice models [Turner2018, Serbyn2021]. These phenomena are nowadays routinely probed across platforms with high control, such as two‑dimensional (2D) electron gases, disordered photonic media, waveguide arrays, and ultracold atoms  [Schwartz2007, Lahini2008, Kondov2011, directVis_Ge2024, PhysRevB103104201, White2020].
@@
Despite the progress on these frontiers, the manner in which finite-size effects and continuum disorder conspire to yield coexisting non-ergodic regimes across the spectrum in realistic systems remains elusive. To address this issue, we here investigate a finite 2D continuum Hamiltonian with prototypical disorder of Gaussian potential bumps serving as a proxy for local impurities. Our analysis reveals regimes that separate across the energy spectrum: (i) conventional AL eigenstates dominating the low-energy sector, (ii) genuinely ergodic delocalized states following random-matrix-theory statistics in the sense of the Bohigas-Giannoni-Schmit conjecture [Bohigas_phys.rev.lett_52_1_1984], while (iii) a set of high-energy scarred states exhibiting intensity patterns that deviate from random-wave predictions of the Berry conjecture [Berry_j.phys.a_10_2083_1977]. Representative eigenstates of these classes are shown in Fig. 1 (upper panel and in Supplemental Material (SM) Fig. S1), illustrating a spectrum of behaviors ranging from AL [Fig. 1(a)], to ergodic delocalization [Fig. 1(b)] and varying degrees of scarring [Figs. 1(c) and (d)]. These regimes are distinguished through an analysis based on complementary observables, including spatial profiles, inverse participation ratios (IPR), multifractal spectra, and deviations from universal level statistics.
Model Setting– We consider a 2D potential profile $V_{\mathrm{ext}}(\mathbf{r})$, confined by Dirichlet hard-wall boundaries to a region of size $L\times L$, where $L$ denotes the number of quantum wells along each direction, together with Gaussian impurities $V_{\text{imp}}(\mathbf{r})$. By utilizing the imaginary-time propagation method [LuukkoP.J.J.2013Itpc], we obtain the lowest $\sim 4\times 10^{3}-10^{4}$ eigenstates of the time-independent Schrödinger equation $H\psi_{n}(\mathbf{r})=E_{n}\psi_{n}(\mathbf{r})$ with
@@
with amplitude $A_{i}$, width $\sigma_{i}$, and position $\mathbf{r}_{i}$, which is a standard approach for modeling imperfections or impurities in quantum wells [hirose_phys.Rev.B_63_075301_2001, hirose_phys.rev.B_65_193305_2002]. The bump positions $\mathbf{r}_{i}$ are sampled independently from a uniform distribution, yielding an average density $\rho$ per unit area. The average impurity width is fixed in all simulations at $\langle\sigma\rangle=0.2$. The dimensionless ratio $\langle A\rangle/V_{0}$, where $\langle A\rangle$ is the average bump amplitude, characterizes the overall disorder strength.
@@
Figure 1 (lower panel) illustrates a typical localization–delocalization crossover in the disorder–energy plane for a system with fixed size $L$ and impurity density $\rho$. The colormap shows $\log_{10}(\mathrm{IPR}_{2})$, and the gray vertical dashed-line marks the scaled well-depth energy $V_{0}$. At energies below and near the well-depth, $E_{n}\lesssim V_{0}$, eigenstates are dominated by potential confinement and increasing disorder strength $\langle A\rangle/V_{0}$ disrupts inter-well resonances (see a representative example in Fig.1(a)). This results in a broadening of the energy window populated by strongly AL states. At energies $E_{n}\gtrsim V_{0}$, the IPR landscape exhibits two quantitatively distinct regimes. One consists of very low IPR eigenstates in which scarring is essentially absent [see Fig. 1(b)]. For these states, the localization length exceeds the system size, and the de Broglie wavelength $\lambda(E)\sim 2\pi/\sqrt{2(E-\langle V\rangle)}$, where $\langle V\rangle$ denotes the expectation value of the total potential energy, remains too large for semiclassical trajectories to resolve the underlying periodic structure [Heller_book].
Upon further increasing the energy, de Broglie wavelength becomes sufficiently short to distinguish the periodic confinement. In this regime, variational scarred states emerge as anisotropic outliers embedded within the delocalized background. As the disorder strength increases, both the low- and intermediate- energy windows [i.e., Fig. 1(a) and (b)] broaden, pushing the scar-bearing region to progressively higher energies. In strongly disordered cases, the remaining scars are weakened and occupy a diminishing fraction of the accessible spectrum, consistent with previous observations [KeskiRahkonen2019b, simo].
@@
As discussed in Fig. 1, at energies $E_{n}\gtrsim V_{0}$, the spectrum contains both delocalized and scarred eigenstates. In 2D, the mean spacing between adjacent energy levels scales as $\sim L^{-2}$ [Girvin_Yang_2019]. Consequently, a fixed set of $10^{4}$ computed eigenstates spans a progressively narrower absolute energy window as the system size increases (see Fig. S3 in SM). Therefore, larger systems require an increasingly greater number of eigenstates to reach energies at which the periodicity of the underlying geometry becomes dynamically relevant. The fixed sampled spectrum is thus dominated by delocalized states, while the scar-bearing window are pushed to higher relative energies.
The coexistence of AL, delocalized, and scarred states within the same sample can thus be explained by the interplay between the energy-dependent $\xi(E)$ and the finite system size. We find that scarred states emerge when (i) $\xi(E)$ becomes comparable to or exceeds $L$, and (ii) de Broglie wavelength becomes sufficiently small to allow eigenfunctions to resolve the underlying PO structure imposed by the bounded geometry. In this regime, the system admits families of unstable and marginally stable POs embedded within the chaotic sea, and interference along these trajectories imprints the characteristic directional, variational scarred patterns onto the corresponding eigenstates [kaplan_ann.phys_264_171_1998, Heller_book].
@@
Figure 3: Level-spacing statistics. (a) Distribution of consecutive level-spacing ratios for disordered systems of sizes $L=4$ and $L=13$. The corresponding symmetrized gap-ratio averages are $\langle\tilde{s}\rangle\approx 0.389$ and $0.524$, respectively. (b) Comparison of weak and strong disorder strengths, $\langle A\rangle/V_{0}=0.1$ and $2$, for $L=5$, yielding $\langle\tilde{s}\rangle\approx 0.388$ and $0.525$, respectively. For reference, the theoretical values are $\langle\tilde{s}\rangle_{\mathrm{Poisson}}\approx 0.386$ and $\langle\tilde{s}\rangle_{\mathrm{GOE}}\approx 0.536$.
@@
Figure S3: $\text{IPR}_{2}$ versus energy for different system sizes, showing progressively narrower absolute energy window as the system size increases .